# Universal Maximum Likelihood (List) Decoding via Fast Vector-Matrix Multiplication

### v1

@@ II-B Extension to Soft Decoding and List Decoding
Observe that the output alphabet $\mathcal{Y}$ and the (conditional) transition probabilities can be arbitrary. Particularly, if the received vector $\boldsymbol{y}$ is considered prior to demodulation—for instance, as the real-valued output of a matched filter—then the output alphabet becomes continuous, i.e., $\mathcal{Y}=\mathbb{R}$. This scenario corresponds to soft decoding, where the channel transition probability $P_{Y|X}$ becomes a conditional probability density function $p_{Y|X}$. It is well known—see, for example, [39, Ch. 4]—that ML decoding in this setting is
$\boldsymbol{c}^{*}$ $=\arg\max_{\boldsymbol{c}\,\in\,\mathscr{C}}p_{Y|X}(\boldsymbol{y}\mid\boldsymbol{c})=\arg\max_{\boldsymbol{c}\in\mathscr{C}}\log\left(\prod_{i=1}^{n}p(y_{i}\mid c_{i})\right)$
$=\arg\max_{\boldsymbol{c}\in\mathscr{C}}\sum_{i=1}^{n}\log p(y_{i}\mid c_{i}).$
@@ II-D Application to Syndrome Decoding for Linear Codes
Combining this with $\mathcal{O}(2^{k})$—the complexity established earlier in the general case of ML decoding of $q$-ary codes over DMC, the overall worst-case complexity for ML decoding of linear codes on the BSC via this approach is $\mathcal{O}(2^{\min\{k,\,n-k\}})$. It is straightforward to extend the method to ML decoding of $q$-ary symmetric channels using $q$-ary linear codes, in which case the total complexity is $\mathcal{O}(q^{\min\{k,\,n-k\}})$.
@@ II-E Extension to Intersymbol-Interference (ISI) Channels
III Conclusion
@@
Vector–matrix multiplication is a common, basic computational routine in numerical linear algebra, optimization, machine learning, and large-scale graph analytics, among many other applications [41]. Its computational complexity remains a central problem in theoretical computer science, where even mild asymptotic improvements can have a wide impact [42]. For a general $m\times n$ matrix $\mathbf{A}$ (where $m<n$) and an $n$-dimensional vector $\boldsymbol{x}$, the straightforward computation of $\mathbf{A}\boldsymbol{x}$ requires $\mathcal{O}(mn)$ time. The key to achieving faster computation lies in exploiting the matrix’s discrete structure. Specifically, when $\mathbf{A}$ has entries drawn from a small finite alphabet—as is the case in our construction—the Mailman algorithm can precompute and reuse partial inner products corresponding to recurring column patterns. This method reduces redundant multiplications and yields a logarithmic improvement, lowering the complexity to $\mathcal{O}(mn/\log n)$ after a one-time preprocessing step. Further refinements improve it to $\mathcal{O}(mn/\log^{2}n)$ [43]. In our setting, this results in an overall runtime of $\mathcal{O}\!\left(\tfrac{q^{k}}{k}\right)$. These finite-alphabet acceleration techniques have also been used recently to speed up the decoding of Hidden Markov Models (HMMs), achieving a logarithmic factor improvement over the Viterbi algorithm [44]. Beyond such logarithmic gains, further improvements appear unlikely: achieving more than a polylogarithmic speedup over Viterbi has been shown to be theoretically hard [45].
Despite these advances, a significant barrier is posed by the widely believed Online Matrix–Vector (OMv) conjecture from theoretical computer science. It asserts that a truly sub-$mn$ time per multiplication—that is, time $o(mn)$ asymptotically smaller than the product of the dimensions—is impossible for general cases. This conjecture is frequently used to establish conditional lower bounds for various computational algorithms across multiple domains [46]. Moreover, additional results show that even with unlimited preprocessing and memory, certain variants of the problem still require near-quadratic time, reinforcing the belief that polynomial improvements are unlikely [47]. For this reason, we consider it improbable that any exact, universal ML decoding method could achieve substantially lower complexity than our proposed algorithm.
@@ Acknowledgment
This was supported in part by NSF-BSF grant FET-2120262. We thank Swapnil Saha, Michael Schleppy, and Karthik C. S. for helpful discussions.

### v2

@@ II-B Extension to Soft Decoding and List Decoding
Observe that the output alphabet $\mathcal{Y}$ and the (conditional) transition probabilities can be arbitrary. Particularly, if the received vector $\boldsymbol{y}$ is considered prior to demodulation—for instance, as the real-valued output of a matched filter—then the output alphabet becomes continuous, i.e., $\mathcal{Y}=\mathbb{R}$. This scenario corresponds to soft decoding, where the channel transition probability $P_{Y|X}$ becomes a conditional probability density function $p_{Y|X}$. It is well known—see, for example, [41, Ch. 4]—that ML decoding in this setting is
$\boldsymbol{c}^{*}$ $=\arg\max_{\boldsymbol{c}\,\in\,\mathscr{C}}p_{Y|X}(\boldsymbol{y}\mid\boldsymbol{c})=\arg\max_{\boldsymbol{c}\in\mathscr{C}}\log\left(\prod_{i=1}^{n}p_{Y|X}(y_{i}\mid c_{i})\right)$
$=\arg\max_{\boldsymbol{c}\in\mathscr{C}}\sum_{i=1}^{n}\log p_{Y|X}(y_{i}\mid c_{i}).$
@@ II-D Application to Syndrome Decoding for Linear Codes
Combining this with $\mathcal{O}(2^{k})$—the complexity established earlier in the general case of ML decoding of $q$-ary codes over DMC, the overall complexity for ML decoding of linear codes on the BSC via this approach is $\mathcal{O}(2^{\min\{k,\,n-k\}})$. It is straightforward to extend the method to ML decoding of $q$-ary symmetric channels using $q$-ary linear codes, in which case the total complexity is $\mathcal{O}(q^{\min\{k,\,n-k\}})$.
@@ II-E Extension to Intersymbol-Interference (ISI) Channels
III Simulation
We evaluate the performance of the proposed decoder against exhaustive search and GRAND [14] over a memoryless binary symmetric channel (BSC), where each bit is independently flipped with probability $p$. The BSC is chosen primarily for implementation simplicity. To ensure a fair comparison, no attempt was made to significantly optimize or parallelize any of the decoding algorithms.
The benchmark decoders operate as follows:
  • Exhaustive search: Upon receiving a vector $\boldsymbol{y}$, the decoder scans the entire codebook $\mathscr{C}$ and outputs the codeword that minimizes the Hamming distance to $\boldsymbol{y}$.
  • GRAND: The decoder generates candidate error patterns $\boldsymbol{e}\in\{0,1\}^{n}$ in decreasing order of likelihood. Over a memoryless BSC with $p<\tfrac{1}{2}$, this ordering coincides with increasing Hamming weight, with ties broken arbitrarily. Decoding terminates when $\boldsymbol{y}-\boldsymbol{e}$ is a valid codeword, i.e.,
    $\boldsymbol{H}(\boldsymbol{y}-\boldsymbol{e})^{\top}=\boldsymbol{0}_{n-k},$
    where $\boldsymbol{H}$ is the code’s parity check matrix, or when a predefined maximum number of guesses, denoted by max_guess, is reached.
All simulations are conducted using binary Reed–Muller codes of order $r$, denoted by $\mathrm{RM}(r,m)$, with $0\leq r\leq m$. Reed–Muller codes have recently attracted renewed interest due to their capacity-achieving behavior over both the BSC and the BEC [30, 42, 1]. An $\mathrm{RM}(r,m)$ code has block length $n=2^{m}$ and dimension $k=\sum_{i=0}^{r}\binom{m}{i}$, resulting in rate $R=k/n$ [33]. For fixed $m$, increasing $r$ increases the dimension and hence the rate of the code.
Simulation results are reported for multiple block lengths $n$ and channel crossover probabilities $p$. In each experiment, 500 codewords from $\mathrm{RM}(1,m)$ are randomly generated and transmitted over the BSC. Decoding time is measured from the moment the corrupted word is produced until a (maximum-likelihood) decoded codeword is declared. The reported decoding times correspond to averages over the 500 trials and are summarized in Fig. 1.
Comparison with Exhaustive Search.
The first plot considers first-order Reed–Muller codes. Since both exhaustive search and the proposed decoder necessarily evaluate all codewords in the codebook, the code rate does not affect the fairness of this comparison. Moreover, for both methods, the bit-flip probability $p$ does not influence the decoding time, which depends only on the block length $n$. The proposed decoder exhibits a noticeably smaller variance in running time; for a fixed block length, any residual variation is attributable solely to implementation-level effects.
As shown in Fig. 1(a), the proposed method consistently outperforms exhaustive search. This behavior is consistent with the theoretical complexity analysis: while exhaustive search has exponential complexity in the parameter $m$, the proposed decoder achieves an additional speedup by a factor of $n$.
Comparison with GRAND.
The second plot considers second-order Reed–Muller codes, which yield moderate-to-high code rates ($R\geq\tfrac{1}{2}$), corresponding to the regime in which GRAND is typically most effective. In this setting, the bit-flip probability $p$ must remain relatively small, since highly noisy channels require codes with large minimum distance, which high-rate codes cannot provide as a consequence of the Singleton bound.
For this reason, the largest block length considered is $n=32$. When $n=64$ and $p=0.065$, the expected number of bit flips is approximately $64\times 0.065\approx 4$. In such cases, GRAND must examine roughly
$\sum_{i=0}^{4}\binom{64}{i}\approx 6.8\times 10^{5}$
candidate error patterns before identifying the correct one, rendering the algorithm prohibitively inefficient.
The comparison with GRAND is therefore more nuanced. Recall that GRAND sequentially subtracts error patterns in decreasing order of likelihood until a valid codeword is found. Consequently, GRAND performs best when $p$ is small or when the code rate is high. Our simulations show that for smaller block lengths and $p=0.03$, GRAND achieves performance comparable to that of the proposed method, since only a limited number of low-weight error patterns need to be tested. However, when the block length increases to $n\geq 32$ and $p=0.065$, the number of plausible error patterns grows rapidly, leading to a substantial degradation in GRAND’s performance relative to the proposed decoder.
(a) Our proposed method vs. exhaustive search
(b) Our proposed method vs. GRAND
Figure 1: Decoding time performance for our proposed algorithm, exhaustive search, and GRAND in decoding the outputs of binary symmetric channels with various crossover probabilities $p$, using first-order Reed-Muller codes $\text{RM}(1,m)$.
IV Conclusion
@@
Vector–matrix multiplication is a common, basic computational routine in numerical linear algebra, optimization, machine learning, and large-scale graph analytics, among many other applications [7]. Its computational complexity remains a central problem in theoretical computer science, where even mild asymptotic improvements can have a wide impact [2]. For a general $m\times n$ matrix $\mathbf{A}$ (where $m<n$) and an $n$-dimensional vector $\boldsymbol{x}$, the straightforward computation of $\mathbf{A}\boldsymbol{x}$ requires $\mathcal{O}(mn)$ time. The key to achieving faster computation lies in exploiting the matrix’s discrete structure. Specifically, when $\mathbf{A}$ has entries drawn from a small finite alphabet—as is the case in our construction—the Mailman algorithm can precompute and reuse partial inner products corresponding to recurring column patterns. This method reduces redundant multiplications and yields a logarithmic improvement, lowering the complexity to $\mathcal{O}(mn/\log n)$ after a one-time preprocessing step. In our setting, this results in a total complexity of $\mathcal{O}(q^{k})$, correspond to a $\Omega(n)$ speedup. This complexity of $\mathcal{O}(q^{k})$ still requires that one examines every possible codeword in the codebook.
Further refinements of the Mailman algorithm improve the complexity of the vector–matrix multiplication routine to $\mathcal{O}(mn/\log^{2}n)$ [48]. In our setting, this yields an overall runtime of $\mathcal{O}\!\left(\tfrac{q^{k}}{k}\right)$. These finite-alphabet acceleration techniques have also been employed to speed up the decoding of Hidden Markov Models (HMMs), achieving a logarithmic-factor improvement over the Viterbi algorithm [8]. Beyond such logarithmic gains, however, further improvements appear unlikely: obtaining more than a polylogarithmic speedup over Viterbi has been shown to be theoretically hard [3], as it would require solving NP-hard problems without exhaustively searching over all possible solutions—an ability that lies at the heart of the P versus NP problem [39].
Moreover, a significant barrier is posed by the widely believed Online Matrix–Vector (OMv) conjecture from theoretical computer science. It asserts that a truly sub-$mn$ time per multiplication—that is, time $o(mn)$ asymptotically smaller than the product of the dimensions—is impossible for general cases. This conjecture is frequently used to establish conditional lower bounds for various computational algorithms across multiple domains [25]. Moreover, additional results show that even with unlimited preprocessing and memory, certain variants of the problem still require near-quadratic time, reinforcing the belief that polynomial improvements are unlikely [22]. For this reason, we consider it hard that any exact, universal ML decoding method could achieve substantially lower complexity than our proposed algorithm.
@@ Acknowledgment
This was supported in part by NSF-BSF grant FET-2120262. We thank Swapnil Saha, Michael Schleppy, Mursalin Habib, and Karthik C. S. for valuable input.